Does the line below pass through point (9,-4)?
x – y = 5

Answer by funmath(2933)

Does the line below pass through point (9,-4)?
x – y = 5
:
(x,y)=(9,-4), let x=9 and y=-4 and see if both sides of the equal sign are equal. If they are, then the point is on the line. If they aren't it isn't.
(9)-(-4)=5
9+4=5
13 not=5
The point is NOT on the line. Therefore the line does NOT pass through the point.
:
For fun, find out that the line does pass through point (9,4).
Happy Calculating!!!